OpenStudy (anonymous):

Three consecutive odd integers are such that the square of the third integer is 9 less than the sum of the squares of the first two. One solution is −3, −1, and 1. Find three other consecutive odd integers that also satisfy the given conditions.

OpenStudy (zpupster):

Let integers are x, x+2, x+4 (x+4)^2 = x^2 + (x+2)^2 - 9 the equation x^2 + 8x + 16 = x^2 + x^2 + 4x + 4 - 9 subtracting the left side of equation 0= x^2 - 4x -21 factor this and test the values

OpenStudy (zpupster):

do you need more help

OpenStudy (mrnood):

(BTW - you know that x=-3 is one of the factors , given in the question)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Is the answer a decimal

OpenStudy (mrnood):

It tells you that they are integers @zpupster has given the correct equation - which is a quadratic: 0= x^2 - 4x -21 You are told that x=-3 is one solution Find the other solution (by factorising) if the other solution is m then the 3 numbers are m, m+2, m+4

OpenStudy (zpupster):

can you factor this ??